OpenStudy (anonymous):

(-5x^3+4)-(-8x^3) math makes no sense to me i need help does anybody know how to do this stuff?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

no one really knows

OpenStudy (anonymous):

40x^6-40x^3

OpenStudy (anonymous):

Thats your answer fan and medal and @ me for more help @cmnorris

OpenStudy (misty1212):

but what you want to do is distribute the minus sign first \[-5x^3+4+8x^3\] is how you start

OpenStudy (misty1212):

then combine like terms

OpenStudy (misty1212):

mister nuclear bird is giving many wrong answers today, beware

OpenStudy (misty1212):

@IQ250 check your work, if you are interested in medals only, you should at least make sure your answer is correct

OpenStudy (adi3):

at first u r gona do this -5x^2 = 25x 25x + 4 = 29

OpenStudy (adi3):

-8x^3 =

OpenStudy (adi3):

@IQ250

OpenStudy (adi3):

@misty1212

OpenStudy (anonymous):

Yes -24X^3

OpenStudy (adi3):

(29x) - (512x) = 29x * -512x =

OpenStudy (adi3):

no I think first you need to squar the number. Thats wht my teacher told me

OpenStudy (adi3):

@IQ250

OpenStudy (adi3):

be quick

OpenStudy (adi3):

anyone reply in 5 minutes

OpenStudy (anonymous):

u cannot square -5x^3

OpenStudy (anonymous):

3x^3+4 is the answers

OpenStudy (adi3):